# Limited-Range Multilinear Off-Diagonal Extrapolation and Weighted Transference Principle

### v1

@@ 2 Limited-Range Multilinear Off-Diagonal Extrapolation
Let $\vec{f}\in\prod_{j=1}^{m}S_{j}^{-1}(L^{p_{1j}}_{{w}_{1j}}(\Omega))$ and $\kappa\in(1,\infty)$. Then $T\vec{f}$ is well defined by Corollary 2.5. Choose $\frac{1}{\lambda}\in(0,\infty)$ sufficiently large such that $\frac{1}{\lambda}>\frac{1}{q_{0}}+\frac{m}{|\gamma_{j}|}$ for all $j\in J$. Then $\tilde{q}_{0},\tilde{q}_{1}\in(1,\infty]$ for $\tilde{q}_{0}:=q_{0}/\lambda$ and $\tilde{q}_{1}:=q_{1}/\lambda$. For $j\in J$ we define $u_{0j}:=m\lambda\tilde{q}_{0}^{\prime}$ and $u_{1j}\in(0,\infty]$ via $\frac{1}{u_{0j}}-\frac{1}{u_{1j}}=\frac{1}{\gamma_{j}}$, so that $\sum_{j=1}^{m}\frac{\lambda}{u_{1j}}=\frac{1}{\tilde{q}_{1}^{\prime}}$ by (8). Thus for $h{v}_{1}^{-\lambda}\in L^{\tilde{q}_{1}^{\prime}}(\Omega)$ with unit norm, the choice $h_{j}:=v_{1j}(|h|v_{1}^{-\lambda})^{\tilde{q}_{1}^{\prime}/u_{1j}}$ yields $0\leq h_{j}{v}_{1j}^{-1}\in L^{u_{1j}}(\Omega)$ with unit norm and $|h|=\prod_{j=1}^{m}h_{j}^{\lambda}$ due to ${v}_{1}:=\prod_{j=1}^{m}{v}_{1j}$.
For $j\in J$, apply Lemma 2.3 with the set of parameters $(u_{0},p_{0},s_{0},r_{0},u,p,s,r)$ replaced by the set $(u_{0j},p_{0j},s_{0j},r_{0j},u_{1j},p_{1j},s_{1j},r_{1j})$ to the weights ${w}_{1j}$, ${v}_{1j}$ and functions $S_{j}f_{j}$ and $h_{j}$. This yields weights $(\vec{{w}}_{0},\vec{{v}}_{0})\in\vec{W}$ such that $\vec{{w}}_{0},\vec{{v}}_{0}]_{(\vec{s}_{0},\vec{r}_{0})}\leq\vec{C}_{\kappa}$ componentwise, and by Hölder’s inequality (recall ${v}_{0}:=\prod_{j\in J}{v}_{0j}$)
$\Big|\int_{\Omega}|T\vec{f}|^{\lambda}h\,\mathrm{d}\mu\Big|$ $\leq\||T\vec{f}|^{\lambda}{v}_{0}^{\lambda}\|_{L^{\tilde{q}_{0}}}\|h{v}_{0}^{-\lambda}\|_{L^{\tilde{q}_{0}^{\prime}}}=\|T\vec{f}\|_{L^{q_{0}}_{{v}_{0}}}^{\lambda}\|h{v}_{0}^{-\lambda}\|_{L^{\tilde{q}_{0}^{\prime}}}$
$\leq\phi([\vec{{w}}_{0},\vec{{v}}_{0}]_{(\vec{s}_{0},\vec{r}_{0})})^{\lambda}\prod_{j=1}^{m}\|S_{j}f_{j}\|_{L^{p_{0j}}_{{w}_{0j}}}^{\lambda}\|h_{j}{v}_{0j}^{-1}\|_{L^{u_{0j}}}^{\lambda}$
$\leq\kappa^{\lambda\sum_{j\in J}\frac{t_{1j}}{|\gamma_{j}|}}\phi(\vec{C})^{\lambda}\prod_{j\in J}\|S_{j}f_{j}\|_{L^{p_{1j}}_{{w}_{1j}}}^{\lambda}.$
Since $\|T\vec{f}\|_{L^{q_{1}}_{{v}_{1}}}^{\lambda}=\||T\vec{f}|^{\lambda}{v}_{1}^{\lambda}\|_{L^{\tilde{q}_{1}}}$ is attained by taking the supremum of $\left|\int_{\Omega}|Tf|^{\lambda}h\,\mathrm{d}\mu\right|$ over all $h{v}_{1}^{-\lambda}\in L^{\tilde{q}_{1}^{\prime}}(\Omega)$ with unit norm, we obtain (10). ∎
@@ 3 Limited-Range Multilinear Off-Diagonal Mixed-Norm Extrapolation
$\|f\|_{L^{\mathbf{p}}(\Omega;X)}:=\left(\int_{\Omega_{1}}\cdots\left(\int_{\Omega_{\ell}}\|f\|_{X}^{p_{1}}\mathrm{d}\mu_{1}\right)^{\frac{p_{2}}{p_{1}}}\cdots\mathrm{d}\mu_{1}\right)^{\frac{1}{p_{\ell}}}<\infty,$
@@
$(\mathbf{t}_{0},\mathbf{t},\boldsymbol{{w}}_{\mathbf{t}},\boldsymbol{{v}}_{\mathbf{t}},\mathbf{a},W)$
@@
and with $a_{i}:=\prod_{k=1}^{i}t_{k}/t_{0k}$
$W:=\begin{cases}\{({w},{v})\}&\text{if }\frac{1}{\boldsymbol{\gamma}}=\mathbf{0},\\ \{(\tilde{w},\tilde{v})\ \lvert\ \tilde{w}\tilde{v}^{-1}=\otimes_{i\in I}({w}_{i}{v}_{i}^{-1})^{a_{i}}\text{ and }[\tilde{w},\tilde{v}]_{(s_{0},r_{0})}<\infty\}&\text{else.}\end{cases}$ (14)
@@
The exponents $a_{i}$ in Definition 3.2 simplify if $s_{\ell}<\ldots<s_{1}<s_{0}$. Indeed, in this situation one has $t_{i}/t_{0i}=s_{i}/s_{i-1}$ for all $i\in I$, and thus $a_{i}=s_{i}/s_{0}$. Similarly, $a_{i}=r_{i}/r_{0}$ if $s_{\ell}>\ldots>s_{1}>s_{0}$.
@@
If ${\mathcal{M}}_{{\mathcal{U}}_{\Lambda_{i}}}$ is bounded from $L^{t_{i}}_{{w}_{t_{i}}}(\Omega_{\Lambda_{i}})$ to $L^{t_{i}}_{{v}_{t_{i}}}(\Omega_{\Lambda_{i}})$ for all $i\in I$, then it holds
$L^{\mathbf{p}}_{\boldsymbol{{w}}}(\Omega)\subseteq\bigcup_{({w}_{0},{v}_{0})\in W}L^{p_{0}}_{{w}_{0}}(\Omega),$
@@
For $\ell=1$, the assertion is simply Corollary 2.5. Let the result be true for all $\ell^{\prime}<\ell$, and consider $f\in L^{\mathbf{p}}_{\boldsymbol{{w}}}(\Omega)$. Define $\tilde{f}\in L^{p_{\ell}}_{{w}_{\ell}}(\Omega_{\ell})$ by $\tilde{f}(x_{\ell}):=\|f(\cdot,x_{\ell})\|_{L^{\tilde{\mathbf{p}}}_{\tilde{\boldsymbol{{w}}}}(\tilde{\Omega})}$, where $\tilde{\mathbf{p}}=(p_{1},\ldots,p_{\ell-1})$, $\tilde{\boldsymbol{{w}}}=({w}_{1},\ldots,{w}_{\ell-1})$, and $\tilde{\Omega}:=\prod_{i=1}^{\ell-1}\Omega_{i}$. The induction hypothesis yields weights $\mu$ and $\nu$ on $\Omega_{\ell}$ with $(\mu,\nu)=({w}_{\ell},{v}_{\ell})$ if $\frac{1}{\gamma_{\ell}}=0$ and $\mu\nu^{-1}=({w}_{\ell}{v}_{\ell}^{-1})^{t_{\ell}/\tilde{t}_{\ell}}$ otherwise, such that $\tilde{f}\in L^{p_{\ell-1}}_{\mu}(\Omega_{\ell})$, i.e. $f\in L^{\tilde{\mathbf{p}}}_{(w_{1},\ldots,w_{\ell-1}\otimes\mu)}(\Omega)$. Thus, again by induction hypothesis, there are weights ${w}_{0}$ and ${v}_{0}$ on $\Omega$ with $({w}_{0},{v}_{0})=(\otimes_{i\in I}{w}_{i},\otimes_{i\in I}{v}_{i})$ if $\frac{1}{\boldsymbol{\gamma}}=0$ and $\frac{{w}_{0}}{{v}_{0}}=(\frac{{w}_{1}}{{v}_{1}})^{a_{1}}\otimes\cdots\otimes(\frac{{w}_{\ell-1}\otimes\mu}{{v}_{\ell-1}\otimes\nu})^{a_{\ell-1}}$ otherwise, such that $f\in L^{p_{0}}_{{w}_{0}}(\Omega)$. Since $(\frac{{w}_{\ell-1}\otimes\mu}{{v}_{\ell-1}\otimes\nu})^{a_{\ell-1}}=(\frac{{w}_{\ell-1}}{{v}_{\ell-1}})^{a_{\ell-1}}\otimes(\frac{{w}_{\ell}}{{v}_{\ell}})^{a_{\ell}}$ by the definition of $\mu$, $\nu$ and $\mathbf{a}$, the result follows. ∎
@@
    $\|{\mathcal{M}}_{{\mathcal{U}}_{\Lambda_{i}}}\|_{L^{t_{ij}}_{{w}_{t_{ij}}}(\Omega_{\Lambda_{i}})\to L^{t_{ij}}_{{v}_{t_{ij}}}(\Omega_{\Lambda_{i}})}\leq\psi_{ij}([{w}_{ij},{v}_{ij}]_{(s_{ij},r_{ij})}),$
@@
    $T:\prod_{j\in J}\bigcup_{({w}_{0j},{v}_{0j})\in W_{j}}S_{j}^{-1}(L^{p_{0j}}_{{w}_{0j}}(\Omega))\to L^{0}(\Omega),$
    and a function $\phi:\mathbb{R}^{m}\to\mathbb{R}$ increasing in each component such that for all $(\vec{{w}}_{0},\vec{{v}}_{0})\in\vec{W}:=W_{1}\times\cdots\times W_{m}$ and $\vec{f}\in\prod_{j\in J}S_{j}^{-1}(L^{\vec{p}_{0}}_{\vec{{w}}_{0}}(\Omega))$ it holds
    $\|T\vec{f}\|_{L^{q_{0}}_{{v}_{0}}(\Omega)}\leq\phi([\vec{{w}}_{0},\vec{{v}}_{0}]_{(\vec{s}_{0},\vec{r}_{0})})\prod_{j\in J}\|S_{j}f_{j}\|_{L^{p_{0j}}_{{w}_{0j}}(\Omega)},$ (16)
Then there is a function $\phi_{\vec{s_{0}},\vec{r_{0}},\vec{\boldsymbol{\gamma}}}:\mathbb{R}^{\ell\times m}\to\mathbb{R}$ increasing in each component such that $T\vec{f}$ is well-defined for all $\vec{f}\in\prod_{j\in J}S_{j}^{-1}(L^{\mathbf{p}_{j}}_{\boldsymbol{{w}}_{j}}(\Omega))$ and
$\|T\vec{f}\|_{L^{\mathbf{q}}_{\boldsymbol{{v}}}(\Omega)}\leq\phi_{\vec{s_{0}},\vec{r_{0}},\vec{\boldsymbol{\gamma}}}([\vec{\boldsymbol{{w}}},\vec{\boldsymbol{{v}}}]_{\vec{\mathbf{s}},\vec{\mathbf{r}}})\prod_{j\in J}\|S_{j}f_{j}\|_{L^{\mathbf{p}_{j}}_{\boldsymbol{{w}}_{j}}(\Omega)}.$ (17)
@@
We establish the conclusion by induction $\ell\in\mathbb{N}$ in a slightly stronger form: There are $\phi_{j}=\phi_{j,\vec{s_{0}},\vec{r_{0}},\vec{\boldsymbol{\gamma}}}:(1,\infty)\times\mathbb{R}^{\ell}\to\mathbb{R}$ increasing in each component with the following property: $T\vec{f}$ is well-defined for all $\vec{f}\in\prod_{j\in J}S_{j}^{-1}(L^{\mathbf{p}_{j}}_{\boldsymbol{{w}}_{j}}(\Omega))$, and for all $\kappa\in(1,\infty)$ it holds
$\|T\vec{f}\|_{L^{\mathbf{q}}_{\boldsymbol{{v}}}(\Omega)}\leq\kappa^{\ell/\alpha}\phi((\phi_{j}(\kappa^{\prime},[\boldsymbol{{w}}_{j},\boldsymbol{{v}}_{j}]_{\mathbf{s}_{j},\mathbf{r}_{j}}))_{j\in J})\prod_{j\in J}\|S_{j}f_{j}\|_{L^{\mathbf{p}_{j}}_{\boldsymbol{{w}}_{j}}(\Omega)}.$ (18)
Let $\vec{f}\in\prod_{j=1}^{m}S_{j}^{-1}(L^{\mathbf{p}_{j}}_{\boldsymbol{{w}}_{j}}(\Omega))$. Then $T\vec{f}$ is well-defined by Lemma 3.3. Fix $\kappa\in(1,\infty)$. Observe that the case $\ell=1$ holds by Theorem 2.6 and the assumption in (i). Suppose that the assertion is true for all $\ell^{\prime}<\ell$. Let $\vec{\mu},\vec{\nu}$ be weight tuples on $\Omega_{\Lambda_{2}}=\prod_{i=2}^{\ell}\Omega_{i}$ with $(\mu_{j},\nu_{j})\in\widetilde{W}_{j}:=W(s_{1j},r_{1j},\otimes_{i=2}^{\ell}({w}_{ij}/{v}_{ij})^{\tilde{a}_{ij}})$, where $\tilde{a}_{ij}:=\prod_{k=2}^{i}t_{kj}/\tilde{t}_{kj}$ Then $\vec{{w}}_{1}\otimes\vec{\mu},\vec{{v}}_{1}\otimes\vec{\nu}]_{(\vec{s}_{1},\vec{r}_{1})}\leq[\vec{{w}}_{1},\vec{{v}}_{1}]_{(\vec{s}_{1},\vec{r}_{1})}[\vec{\mu},\vec{\nu}]_{(\vec{s}_{1},\vec{r}_{1})}<\infty$. Estimate (16) is valid by assumption for all $(\vec{{w}}_{0},\vec{{v}}_{0})$ with $({w}_{0j},{v}_{0j})\in W(s_{0j},r_{0j},(\frac{{w}_{1j}\otimes\mu_{j}}{{v}_{1j}\otimes\nu_{j}})^{t_{1j}/t_{01j}})\subseteq W_{j}$ for all $j\in J$ and all $\vec{g}\in\prod_{j\in J}S_{j}^{-1}(L^{\vec{p}_{0}}_{\vec{{w}}_{0}}(\Omega))$, and serves as input (9) to Theorem 2.6. This yields that $T\vec{g}$ is well-defined for all $\vec{g}\in\prod_{j\in J}S_{j}^{-1}(L^{p_{1j}}_{{w}_{1j}\otimes\mu_{j}}(\Omega))$ and
$\|T\vec{g}\|_{L^{q_{1}}_{{v}_{1}\otimes\nu}}\leq\kappa^{\frac{1}{\alpha}}\phi((C_{1j}(\kappa^{\prime},[{w}_{1j},{v}_{1j}]_{(s_{1j},r_{1j})},[\mu_{j},\nu_{j}]_{(s_{1j},r_{1j})}))_{j\in J})\prod_{j\in J}\|S_{j}g_{j}\|_{L^{p_{1j}}_{{w}_{1j}\otimes\mu_{j}}},$ (19)
where $C_{1j}(\rho,x_{j},y_{j}):=(\rho\psi_{1j}(x_{j}y_{j}))^{t_{1j}/|\gamma_{1j}|}(x_{j}y_{j})^{t_{1j}/t_{01j}}$. Let $\widetilde{S}_{j}:V_{j}\to L^{0}(\Omega_{\Lambda_{2}})$ be given by $\widetilde{S}_{j}g_{j}(x^{\prime}):=\|S_{j}g_{j}(x^{\prime},\cdot)\|_{L^{p_{1j}}_{{w}_{1j}}(\Omega_{1})}$ and
$\widetilde{T}:\prod_{j=1}^{m}\bigcup_{({w}_{0j},{v}_{0j})\in\widetilde{W}_{j}}\widetilde{S}_{j}^{-1}(L^{p_{0j}}_{{w}_{0j}}(\Omega_{\Lambda_{2}}))\to L^{0}(\Omega_{\Lambda_{2}}),\qquad\widetilde{T}\vec{g}(x^{\prime}):=\|T\vec{g}(x^{\prime},\cdot)\|_{L^{q_{1}}_{{w}_{1}}(\Omega_{1})},$
@@
$\|\widetilde{T}\vec{g}\|_{L^{q_{1}}_{\nu}(\Omega_{\Lambda_{2}})}\leq\tilde{\phi}([\vec{\mu},\vec{\nu}]_{(\vec{s}_{1},\vec{r}_{1})})\prod_{j\in J}\|\widetilde{S}_{j}g_{j}\|_{L^{p_{1j}}_{\mu_{j}}(\Omega_{\Lambda_{2}})},$

### v2

@@ 2 Limited-Range Multilinear Off-Diagonal Extrapolation
Let $\vec{f}\in\prod_{j=1}^{m}S_{j}^{-1}(L^{p_{1j}}_{{w}_{1j}}(\Omega))$ and $\kappa\in(1,\infty)$. Then $T\vec{f}$ is well defined by Corollary 2.5. Choose $\frac{1}{\lambda}\in(0,\infty)$ sufficiently large such that $\frac{1}{\lambda}>\frac{1}{q_{0}}+\frac{m}{|\gamma_{j}|}$ for all $j\in J$. Then $\tilde{q}_{0},\tilde{q}_{1}\in(1,\infty]$ for $\tilde{q}_{0}:=q_{0}/\lambda$ and $\tilde{q}_{1}:=q_{1}/\lambda$. For $j\in J$ we define $u_{0j}:=m\lambda\tilde{q}_{0}^{\prime}$ and $u_{1j}\in(0,\infty]$ via $\frac{1}{u_{0j}}-\frac{1}{u_{1j}}=\frac{1}{\gamma_{j}}$, so that $\sum_{j=1}^{m}\frac{\lambda}{u_{1j}}=\frac{1}{\tilde{q}_{1}^{\prime}}$ by (8). Thus for $h{w}_{1}^{-\lambda}\in L^{\tilde{q}_{1}^{\prime}}(\Omega)$ with unit norm, the choice $h_{j}:={w}_{1j}(|h|{w}_{1}^{-\lambda})^{\tilde{q}_{1}^{\prime}/u_{1j}}$ yields $0\leq h_{j}{w}_{1j}^{-1}\in L^{u_{1j}}(\Omega)$ with unit norm and $|h|=\prod_{j=1}^{m}h_{j}^{\lambda}$ due to ${w}_{1}:=\prod_{j=1}^{m}{w}_{1j}$.
For $j\in J$, apply Lemma 2.3 with the set of parameters $(u_{0},p_{0},s_{0},r_{0},u,p,s,r)$ replaced by the set $(u_{0j},p_{0j},s_{0j},r_{0j},u_{1j},p_{1j},s_{1j},r_{1j})$ to the weights ${w}_{1j}$, ${v}_{1j}$ and functions $S_{j}f_{j}$ and $h_{j}$. This yields weights $(\vec{{w}}_{0},\vec{{v}}_{0})\in\vec{W}$ such that $\vec{{w}}_{0},\vec{{v}}_{0}]_{(\vec{s}_{0},\vec{r}_{0})}\leq\vec{C}_{\kappa}$ componentwise, and by Hölder’s inequality (recall ${w}_{0}:=\prod_{j\in J}{w}_{0j}$)
$\Big|\int_{\Omega}|T\vec{f}|^{\lambda}h\,\mathrm{d}\mu\Big|$ $\leq\||T\vec{f}|^{\lambda}{w}_{0}^{\lambda}\|_{L^{\tilde{q}_{0}}}\|h{w}_{0}^{-\lambda}\|_{L^{\tilde{q}_{0}^{\prime}}}=\|T\vec{f}\|_{L^{q_{0}}_{{w}_{0}}}^{\lambda}\|h{w}_{0}^{-\lambda}\|_{L^{\tilde{q}_{0}^{\prime}}}$
$\leq\phi([\vec{{w}}_{0},\vec{{v}}_{0}]_{(\vec{s}_{0},\vec{r}_{0})})^{\lambda}\prod_{j=1}^{m}\|S_{j}f_{j}\|_{L^{p_{0j}}_{{v}_{0j}}}^{\lambda}\|h_{j}{w}_{0j}^{-1}\|_{L^{u_{0j}}}^{\lambda}$
$\leq\kappa^{\lambda\sum_{j\in J}\frac{t_{1j}}{|\gamma_{j}|}}\phi(\vec{C})^{\lambda}\prod_{j\in J}\|S_{j}f_{j}\|_{L^{p_{1j}}_{{v}_{1j}}}^{\lambda}.$
Since $\|T\vec{f}\|_{L^{q_{1}}_{{w}_{1}}}^{\lambda}=\||T\vec{f}|^{\lambda}{w}_{1}^{\lambda}\|_{L^{\tilde{q}_{1}}}$ is attained by taking the supremum of $\left|\int_{\Omega}|Tf|^{\lambda}h\,\mathrm{d}\mu\right|$ over all $h{w}_{1}^{-\lambda}\in L^{\tilde{q}_{1}^{\prime}}(\Omega)$ with unit norm, we obtain (10). ∎
@@ 3 Limited-Range Multilinear Off-Diagonal Mixed-Norm Extrapolation
$\|f\|_{L^{\mathbf{p}}(\Omega;X)}:=\left(\int_{\Omega_{\ell}}\cdots\left(\int_{\Omega_{1}}\|f\|_{X}^{p_{1}}\mathrm{d}\mu_{1}\right)^{\frac{p_{2}}{p_{1}}}\cdots\mathrm{d}\mu_{\ell}\right)^{\frac{1}{p_{\ell}}}<\infty,$
@@
$(\mathbf{t}_{0},\mathbf{t},\boldsymbol{{w}}_{\mathbf{t}},\boldsymbol{{v}}_{\mathbf{t}},W)$
@@
and (recall ${w}:=\otimes_{i\in I}{w}_{i}$, ${v}:=\otimes_{i\in I}{v}_{i}$)
$W:=\begin{cases}\{({w},{v})\}&\text{if }\frac{1}{\boldsymbol{\gamma}}=\mathbf{0},\\ \{(\tilde{w},\tilde{v})\ \lvert\ \tilde{w}\tilde{v}^{-1}={w}{v}^{-1}\text{ and }[\tilde{w},\tilde{v}]_{(s_{0},r_{0})}<\infty\}&\text{else.}\end{cases}$ (14)
@@
If ${\mathcal{M}}_{{\mathcal{U}}_{\Lambda_{i}}}$ is bounded from $L^{t_{i}}_{{v}_{t_{i}}}(\Omega_{\Lambda_{i}})$ to $L^{t_{i}}_{{w}_{t_{i}}}(\Omega_{\Lambda_{i}})$ for all $i\in I$, then it holds
$L^{\mathbf{p}}_{\boldsymbol{{v}}}(\Omega)\subseteq\bigcup_{({w}_{0},{v}_{0})\in W}L^{p_{0}}_{{v}_{0}}(\Omega).$
@@
For $\ell=1$, the assertion is simply Corollary 2.5. Let the result be true for all $\ell^{\prime}<\ell$, and consider $f\in L^{\mathbf{p}}_{\boldsymbol{{v}}}(\Omega)$. Define $\tilde{f}\in L^{p_{\ell}}_{{v}_{\ell}}(\Omega_{\ell})$ by $\tilde{f}(x_{\ell}):=\|f(\cdot,x_{\ell})\|_{L^{\tilde{\mathbf{p}}}_{\tilde{\boldsymbol{{v}}}}(\tilde{\Omega})}$, where $\tilde{\mathbf{p}}=(p_{1},\ldots,p_{\ell-1})$, $\tilde{\boldsymbol{{w}}}=({w}_{1},\ldots,{w}_{\ell-1})$, and $\tilde{\Omega}:=\prod_{i=1}^{\ell-1}\Omega_{i}$. The induction hypothesis yields weights $\mu$ and $\nu$ on $\Omega_{\ell}$ with $(\mu,\nu)=({w}_{\ell},{v}_{\ell})$ if $\frac{1}{\gamma_{\ell}}=0$ and $\mu\nu^{-1}={w}_{\ell}{v}_{\ell}^{-1}$ otherwise, such that $\tilde{f}\in L^{p_{\ell-1}}_{\nu}(\Omega_{\ell})$, i.e. $f\in L^{\tilde{\mathbf{p}}}_{({v}_{1},\ldots,{v}_{\ell-1}\otimes\nu)}(\Omega)$. Thus, again by induction hypothesis, there are weights ${w}_{0}$ and ${v}_{0}$ on $\Omega$ with $({w}_{0},{v}_{0})=(\otimes_{i\in I}{w}_{i},\otimes_{i\in I}{v}_{i})$ if $\frac{1}{\boldsymbol{\gamma}}=0$ and $\frac{{w}_{0}}{{v}_{0}}=(\frac{{w}_{1}}{{v}_{1}})\otimes\cdots\otimes(\frac{{w}_{\ell-1}\otimes\mu}{{v}_{\ell-1}\otimes\nu})$ otherwise, such that $f\in L^{p_{0}}_{{v}_{0}}(\Omega)$. Since $(\frac{{w}_{\ell-1}\otimes\mu}{{v}_{\ell-1}\otimes\nu})=(\frac{{w}_{\ell-1}}{{v}_{\ell-1}})\otimes(\frac{{w}_{\ell}}{{v}_{\ell}})$ by the definition of $\mu$ and $\nu$, the result follows. ∎
@@
    $\|{\mathcal{M}}_{{\mathcal{U}}_{\Lambda_{i}}}\|_{L^{t_{ij}}_{{v}_{t_{ij}}}(\Omega_{\Lambda_{i}})\to L^{t_{ij}}_{{w}_{t_{ij}}}(\Omega_{\Lambda_{i}})}\leq\psi_{ij}([{w}_{ij},{v}_{ij}]_{(s_{ij},r_{ij})}),$
@@
    $T:\prod_{j\in J}\bigcup_{({w}_{0j},{v}_{0j})\in W_{j}}S_{j}^{-1}(L^{p_{0j}}_{{v}_{0j}}(\Omega))\to L^{0}(\Omega),$
    and a function $\phi:\mathbb{R}^{m}\to\mathbb{R}$ increasing in each component such that for all $(\vec{{w}}_{0},\vec{{v}}_{0})\in\vec{W}:=W_{1}\times\cdots\times W_{m}$ and $\vec{f}\in\prod_{j\in J}S_{j}^{-1}(L^{\vec{p}_{0}}_{\vec{{v}}_{0}}(\Omega))$ it holds
    $\|T\vec{f}\|_{L^{q_{0}}_{{w}_{0}}(\Omega)}\leq\phi([\vec{{w}}_{0},\vec{{v}}_{0}]_{(\vec{s}_{0},\vec{r}_{0})})\prod_{j\in J}\|S_{j}f_{j}\|_{L^{p_{0j}}_{{v}_{0j}}(\Omega)},$ (16)
Then there is a function $\phi_{\vec{s_{0}},\vec{r_{0}},\vec{\boldsymbol{\gamma}}}:\mathbb{R}^{\ell\times m}\to\mathbb{R}$ increasing in each component such that $T\vec{f}$ is well-defined for all $\vec{f}\in\prod_{j\in J}S_{j}^{-1}(L^{\mathbf{p}_{j}}_{\boldsymbol{{v}}_{j}}(\Omega))$ and
$\|T\vec{f}\|_{L^{\mathbf{q}}_{\boldsymbol{{w}}}(\Omega)}\leq\phi_{\vec{s_{0}},\vec{r_{0}},\vec{\boldsymbol{\gamma}}}([\vec{\boldsymbol{{w}}},\vec{\boldsymbol{{v}}}]_{\vec{\mathbf{s}},\vec{\mathbf{r}}})\prod_{j\in J}\|S_{j}f_{j}\|_{L^{\mathbf{p}_{j}}_{\boldsymbol{{v}}_{j}}(\Omega)}.$ (17)
@@
We establish the conclusion by induction $\ell\in\mathbb{N}$ in a slightly stronger form: There are $\phi_{j}=\phi_{j,\vec{s_{0}},\vec{r_{0}},\vec{\boldsymbol{\gamma}}}:(1,\infty)\times\mathbb{R}^{\ell}\to\mathbb{R}$ increasing in each component with the following property: $T\vec{f}$ is well-defined for all $\vec{f}\in\prod_{j\in J}S_{j}^{-1}(L^{\mathbf{p}_{j}}_{\boldsymbol{{v}}_{j}}(\Omega))$, and for all $\kappa\in(1,\infty)$ it holds
$\|T\vec{f}\|_{L^{\mathbf{q}}_{\boldsymbol{{w}}}(\Omega)}\leq\kappa^{\ell/\alpha}\phi((\phi_{j}(\kappa^{\prime},[\boldsymbol{{w}}_{j},\boldsymbol{{v}}_{j}]_{\mathbf{s}_{j},\mathbf{r}_{j}}))_{j\in J})\prod_{j\in J}\|S_{j}f_{j}\|_{L^{\mathbf{p}_{j}}_{\boldsymbol{{v}}_{j}}(\Omega)}.$ (18)
Let $\vec{f}\in\prod_{j=1}^{m}S_{j}^{-1}(L^{\mathbf{p}_{j}}_{\boldsymbol{{v}}_{j}}(\Omega))$. Then $T\vec{f}$ is well-defined by Lemma 3.3. Fix $\kappa\in(1,\infty)$. Observe that the case $\ell=1$ holds by Theorem 2.6 and the assumption in (i). Suppose that the assertion is true for all $\ell^{\prime}<\ell$. Let $\vec{\mu},\vec{\nu}$ be weight tuples on $\Omega_{\Lambda_{2}}=\prod_{i=2}^{\ell}\Omega_{i}$ with $(\mu_{j},\nu_{j})\in\widetilde{W}_{j}:=W(s_{1j},r_{1j},\otimes_{i=2}^{\ell}{w}_{ij}{v}_{ij}^{-1})$. Then $\vec{{w}}_{1}\otimes\vec{\mu},\vec{{v}}_{1}\otimes\vec{\nu}]_{(\vec{s}_{1},\vec{r}_{1})}\leq[\vec{{w}}_{1},\vec{{v}}_{1}]_{(\vec{s}_{1},\vec{r}_{1})}[\vec{\mu},\vec{\nu}]_{(\vec{s}_{1},\vec{r}_{1})}<\infty$. Estimate (16) is valid by assumption for all $(\vec{{w}}_{0},\vec{{v}}_{0})$ with $({w}_{0j},{v}_{0j})\in W(s_{0j},r_{0j},\frac{{w}_{1j}\otimes\mu_{j}}{{v}_{1j}\otimes\nu_{j}})\subseteq W_{j}$ for all $j\in J$ and all $\vec{g}\in\prod_{j\in J}S_{j}^{-1}(L^{\vec{p}_{0}}_{\vec{{v}}_{0}}(\Omega))$, and serves as input (9) to Theorem 2.6. This yields that $T\vec{g}$ is well-defined for all $\vec{g}\in\prod_{j\in J}S_{j}^{-1}(L^{p_{1j}}_{{v}_{1j}\otimes\mu_{j}}(\Omega))$ and
$\|T\vec{g}\|_{L^{q_{1}}_{{w}_{1}\otimes\mu}}\leq\kappa^{\frac{1}{\alpha}}\phi((C_{1j}(\kappa^{\prime},[{w}_{1j},{v}_{1j}]_{(s_{1j},r_{1j})},[\mu_{j},\nu_{j}]_{(s_{1j},r_{1j})}))_{j\in J})\prod_{j\in J}\|S_{j}g_{j}\|_{L^{p_{1j}}_{{v}_{1j}\otimes\nu_{j}}},$ (19)
where $C_{1j}(\rho,x_{j},y_{j}):=(\rho\psi_{1j}(x_{j}y_{j}))^{t_{1j}/|\gamma_{1j}|}(x_{j}y_{j})^{t_{1j}/t_{01j}}$. Let $\widetilde{S}_{j}:V_{j}\to L^{0}(\Omega_{\Lambda_{2}})$ be given by $\widetilde{S}_{j}g_{j}(x^{\prime}):=\|S_{j}g_{j}(x^{\prime},\cdot)\|_{L^{p_{1j}}_{{v}_{1j}}(\Omega_{1})}$ and
$\widetilde{T}:\prod_{j=1}^{m}\bigcup_{({w}_{0j},{v}_{0j})\in\widetilde{W}_{j}}\widetilde{S}_{j}^{-1}(L^{p_{0j}}_{{v}_{0j}}(\Omega_{\Lambda_{2}}))\to L^{0}(\Omega_{\Lambda_{2}}),\qquad\widetilde{T}\vec{g}(x^{\prime}):=\|T\vec{g}(x^{\prime},\cdot)\|_{L^{q_{1}}_{{w}_{1}}(\Omega_{1})},$
@@
$\|\widetilde{T}\vec{g}\|_{L^{q_{1}}_{\mu}(\Omega_{\Lambda_{2}})}\leq\tilde{\phi}([\vec{\mu},\vec{\nu}]_{(\vec{s}_{1},\vec{r}_{1})})\prod_{j\in J}\|\widetilde{S}_{j}g_{j}\|_{L^{p_{1j}}_{\nu_{j}}(\Omega_{\Lambda_{2}})},$